# Loss Functions for Measuring the Accuracy of Nonnegative Cross-Sectional Predictions

### v2

@@ Abstract
Measuring the accuracy of cross-sectional predictions is a subjective problem. Generally, this problem is avoided. In contrast, this paper confronts subjectivity up front by eliciting an impartial decision-maker’s preferences. These preferences are embedded into an axiomatically-derived loss function, the simplest version of which is described. The parameters of the loss function can be estimated by linear regression. Specification tests for this function are described. This framework is extended to weighted averages of estimates to find the optimal weightings. Rescalings to account for changes in control data or base year data are considered. A special case occurs when the predictions represent resource allocations: the apportionment literature is used to construct the Webster-Saint Lagüe Rule, a particular parametrization of the loss function. These loss functions are compared to those existing in the literature. Finally, a bias measure is created that uses signed versions of these loss functions.
@@ 1 Introduction
The accuracy of cross-sectional predictions is of importance to a large number of their users: governments, investors, and so on. These data include, for example, population, employment or income per capita by geographic area. The outstanding characteristics of cross-sectional data are a great range of values and systematically varying error variances and coefficients of variation. Predictions encompass both forecasts and estimates, the only difference, generally irrelevant to this article, is the date being predicted relative to the date the predictions are made. In the context of predictions, accuracy is generally a subjective concept. [^1] Compare Lindley (1953, p. 46): “…I feel that loss is not easily separated in our minds from beliefs.” Later we develop loss functions to reflect beliefs. Even if objective measures of scale can be used, they may not necessarily coincide with the user’s concept of loss. This paper axiomatically developing loss functionns, which measure the “badness” of errors as a function of the sizes of the error and the actual values, or, equivalently, the predicted and actual values. This approach is in contrast to the common use of measures based on time series. [^2] Armstrong and Collopy (1992) and Fildes (1992), to give two examples, are quite clear about the time series basis of the error measures they use. These measures were developed to measure variations in the level of a single variable over time and are, thus, inapplicable to cross-sectional data. [^3] NRC (1980), is one of the few examples to use cross-sectional criteria: “‘To minimize errors in allocations’ is a vague statement which allows several interpretations (NRC, 1980, p.86).” The allocations are revenues directed to states and localities in proportion to their populations. I am indebted to Bashiruddin Ahmed for providing this example. Spencer (1980, 1985, 1986) also uses cross-sectional criteria. Initially, an impartial decision-maker is assumed to exist and asked to specify values of his loss function (disguised) for various combinations of errors and actual values. The decision-maker is assumed to be interested solely in the overall accuracy of the predictions and not to have an interest in the accuracy for a particular area or set of areas. The context of the predictions is also assumed fixed. [^4] I’d like to thank Stan Smith for pointing this out. Linear regression is then used to estimate the parameters of the loss function. [^5] The assumption of an impartial decision-maker and the statistical construction of the loss function provide an answer Spencer’s (1986) criticisms about whose preferences are to be incorporated into a von Neumann-Morgenstern loss function and how. These constructions are very much along the lines of Wald (1950) and Lindley (1953, 1985). An upshot of this process is that, in general, no single “ideal” measure of accuracy exists, as the evaluation of accuracy depends on the evaluator. [^6] Tukey (1986, p. 408) writes in the context of cross-sectional population estimates: “‘Optimum’ estimates, for example, are only optimum under narrow specifications that do not hold exactly in practice.” (Original underline) In the special case in which predictions represent resource allocations, the apportionment literature is used to establish the Webster-Saint Lagüe Rule as the appropriate loss function.
@@
Section 2 axiomatically constructs the simple loss function that we will use. It further specifies assumptions and develops the simplest loss function that satisfies those assumptions also has the property that the loss associated with a given absolute relative error [^7] Equivalently, one can also refer to the absolute percentage error (APE), but this will complicate the arguments made later on. APE will appear, however, in the example in Section 6. increases in the actual value. Section 3 discusses estimation of this loss function, along with specification tests. Different sets of predictions may have systematic, offsetting biases. The decision-maker may wish to form weighted averages of these predictions to produce a set of new predictions. Subection 3.4 provides a method to search for the optimal weightings which minimize the total loss function. Section 4 argues for the use of the Webster-Saint Lagüe Rule when predictions can be viewed as apportionments. Given a set of predicted and actual values, the total loss function is the sum of the loss functions for the individual observations. Section 5 provides an example of the use of loss functions, comparing the Webster-Saint Lagüe Rule to the commonly used mean absolute percentage error (MAPE).
@@ 2 The Loss Function
This section describes the assumptions used to generate the loss function, $L(P;A)$, where $P$ is the predicted value and $A$ is the actual value. [^8] Muhsam (1966) was apparently the first to use this form of loss in the context of predicting population. He was concerned with the loss associated with a point forecast of the population of a single geographic area. He assumed continuity of loss and a version of Assumption 2 below. I’d like to thank Jay Siegel for this point. After the assumptions are made, the simplest form of $L$ which satisfies them is specified. Restrictions on the values of the parameters of $L$ which make it increase in $A$ for a given relative error are then specified. The total loss $\mathcal{L}$ is the sum of the losses associated with each observation $i$: $\mathcal{L}=\sum_{i}^{n}L(P_{i}-A_{i})$, where n is the number of observations. [^9] In terms of Spencer (1986), $L$ is a component loss function and $\mathcal{L}$ is the loss function. The U.S. Census Bureau also uses this terminology for evaluating census adjustments. While other forms of $\mathcal{L}$ are possible, the summation corresponds to the concept of additively separable utility and obeys the von Neumann-Morgenstern expected utility axioms. [^10] That is, loss is the negative of utility. Additive losses are equivalent to additive utility. The von Neumann-Morgenstern expected utility axioms imply that the expected utility of a gamble is equal to the utilities of the outcomes multiplied by their probabilities. In terms of loss, the expected loss when losses of values $L_{1}$ and $L_{2}$ occur with probabilities $p_{1}$ and $p_{2}$, respectively, is $p_{1}L_{1}+p_{2}L_{2}$. This approach is similar to Lindley’s (1985, p. 59) commandment to “Go forth and maximise your expected utility.” (Original quotes) Lindley (1985) builds his book around the concept of making decisions that maximize a decision-maker’s expected utility. Since expectation requires specifying probability distributions, which we do not do, our commandment can be stated as “Go forth and maximize utility.”
@@
The simplest function which satisfies Assumptions 1–3 and admits Property 1 below is the Cobb-Douglas function
@@ 3.2 Determining the Loss Function at a Set of Points and Estimating it Globally
where $L_{j}>0,a$ is the constant of integration, an ignorable scaling factor, and $u_{j}$ is the regression residual. The values of $p$ and $q$ found from this regression become the parameters of the loss function. One caveat is in order: if the decision-maker reports a $U_{j}=100$, then and its logarithm becomes $\infty$- ∞. The fix for this is to either delete this observation from the analysis or to choose a small $\delta>0$ to bound the $L_{j}$ from below.
@@ 3.2.1 Multiple Decision Makers
A final alternative is to adopt a loss function which works “well” in most cases as a convention in the manner of Keyfitz (1979). The Mean Absolute Percentage Error (MAPE,) discussed in Section 6, is frequently used in this sense. Another alternative is the Webster-Saint Lagüe Rule, discussed in Section 5. Webster’s Rule is a particular parametrization of the loss functions developed in the present paper. That the chosen loss function need not work well in every case has been explicated by Lerner (1957, p. 441) in the context of legal rules:
@@ 3.3 Specification Tests
While equations (1a) and (1b) are the simplest form that the loss function can take, they are hardly the only forms. In fact, an infinite number of forms of loss functions are possible. One quick specification test is to look at the sum $p+q$. If it is nonpositive, then it violates Property 1. Either the equation is misspecified, [^23] Remember that equations (1a) and (1b) represent only one of an infinite number of loss functions that satisfy Assumptions 1-3 and Property 1. or the decision-maker’s preferences violate Property 1.242424In some contexts, this may be reasonable. The form of the function may be misspecified, for which many tests are available. Heteroscedasticity may be present, which can be accounted for by weighted regression.
@@ 3.4 Optimal Averages of Predictions
Given two or more sets of predictions, it may be possible to improve on their accuracy by finding an optimal weighted average of these predictions. Suppose there are two sets of estimates $\mathbf{P}^{1}$ and $\mathbf{P}^{2}$, with associated total losses $\mathcal{L}^{1}$ and $\mathcal{L}^{2}$. Let $w_{1}$ and $w_{2}$ be weights such that $w_{1}+w_{2}=1$. By varying the values of $w_{1}$ and $w_{2}$, one creates new sets of predictions of the form and computes the total loss associated with each set of weights. [^25] The new estimates may have to be constrained to sum up to a predetermined overall total or several predetermined subtotals. See Subsection 4.1. The total loss minimizing set of estimates and weights is then chosen. The optimal search method is unclear: this probably requires grid-searching. With $k>2$ estimate sets, this becomes a possibly expensive mesh search in $k-1$ dimensions.
@@ 4 Predictions as Apportionments: The Webster-Saint Lagüe Rule
Often predictions represent apportionments, that is, resource or legislative seat allocations. The U.S. decennial census is used to apportion the House of Representatives. Population estimates are used by some states to apportion tax revenues to localities. One would like to have the rule used for evaluating the predictions satisfy various fairness criteria. Balinski and Young (2001) show that the Webster-Saint Lagüe Rule satisfies a large number of fairness criteria. [^29] However, while Arrow’s( 1950) Impossibility Theorem implies that no voting system can simultaneously satisfy all fairness criteria, the Webster-Saint Lagüe Rule is the only apportionment rule considered by Balinski and Young (1979, 2001) and Ernst (1992) that can be represented by minimizing the sum of loss functions satisifying Assumptions 1-3 and Property 1 (Spencer, 1985). I would like to thank Gregg Diffendal for informing me of Ernst (1992). Most of these criteria result from the Webster-Saint Lagüe Rule being a divisor method. the Webster-Saint Lagüe Rule method is the only divisor method which has an admissible (in the present paper) loss function representation: $p=2$ and $q=-1$, (Balinski and Young, 1985) See Subsection 6.5 below for other apportionment-based loss functions. An additional normative criterion of particular interest is unbiasedness: for any two disjoint groups of areas, with all members in one group larger than those in the other, the probability that the method favors large areas equals the probability that it favors the small areas. (Balinksi and Young, 2001, Theorem 5.3)
@@ 4.1 An A Priori Probabilistic Interpretation of the Webster-Saint Lagüe Rule
$=A_{i}\mathrm{Var}\left(\sum_{k=1}^{A_{i}}\eta_{i}^{j}\right)$ (6d)
$=A_{i}\sum_{k=1}^{A_{i}}\mathrm{Var}(\eta_{i}^{j})$ (6e)
@@
where equation (6e) follows from the variance of the sum of independent random variables being equal to the sum of their variances and equation (6f) follows from the definition of $\sigma^{2}_{j}$. Therefore,
@@
Thus, the Webster-Saint Lagüe rule produces estimates of the variances of each observation, $A_{i}\hat{\sigma}_{ij}^{2}$. Using its minimization to select predictions chooses the estimates set with the lowest $\hat{\sigma}_{j}^{2}$. Relaxing the the identical variance assumption so that $\sigma^{2}_{ij}$ is not constant causes this rule to estimate the average variance $\hat{\sigma}^{2}_{j}=\frac{1}{n}\sum_{k=1}^{n}\hat{\sigma}_{kj}^{2}$. Now, the Webster-Saint Lagüe loss minimization selects the predictions set with the lowest average estimated variance. This is robust to small departures from the assumptions.
@@ 5 Example of Evaluating Predictions Using Loss Functions
Three Scenarios are used to compare the results of an evaluation using a loss function to those obtained by using MAPE. Scenario 1 is the baseline scenario with $\mathrm{APE}_{i})\equiv 2$ and $\mathcal{L}_{W}/n=11.08$. In Scenario 2, $\mathrm{APE}_{i}$ is reduced to 1 for $i\leq 5$, but $\mathrm{APE}_{6}i$ increases to 10. That is, all but the smallest areas have their APEs halved, but the very smallest area’s APE increases by a factor of 5. MAPE increases to 2.5, but $\mathcal{L}_{W}/n$ falls to 2.9. Thus, MAPE ranks Scenario 2 as being less accurate than Scenario 1, even though the individual errors are smaller except for the very smallest area. On the other hand, the loss function takes into account the size of the smallest area and discounts its accuracy loss and considers Scenario 2 to be more accurate. In Scenario 3, $\mathrm{APE}_{i}$ falls by 15% to 1.7 for $i=2\leq i\leq 5$, rises by 50% to 3 in the largest area, and is unchanged in the smallest area. MAPE falls slightly to 1.97, but $\mathcal{L}_{W}/n$ rises to 18.19. Thus, MAPE considers Scenario 3 to be superior to Scenario 1, as a result of the general reduction in the $\mathrm{APE}_{i}$, in spite of the major loss in accuracy in the largest area. The loss function, on the other hand, puts a large weight on the accuracy loss in area 1 and increases its error measure relative to Scenario 1. Thus, the loss function puts increasing weight on an error as the size of the area increases. Putting all of these together, we find that MAPE and the Webster-Saint Lagüe loss function produce exactly opposite rankings of the Scenarios.

### v3

@@ Abstract
Measuring the accuracy of cross-sectional predictions is a subjective problem. Generally, this problem is avoided. In contrast, this paper confronts subjectivity up front by eliciting an impartial decision-maker’s preferences. These preferences are embedded into an axiomatically-derived loss function, one of the simplest version of which is described. The parameters of the loss function can be estimated by linear regression. Specification tests for this function are described. This framework is extended to weighted averages of estimates to find the optimal weightings. A special case occurs when the predictions represent resource allocations: the apportionment literature is used to construct the Webster-Saint Lagüe Rule, a particular parametrization of the loss function. These loss functions are compared to those existing in the literature. Finally, a family of bias measures are created using signed versions of these loss functions.
@@ 1 Introduction
The accuracy of cross-sectional predictions is of importance to a large number of their users: governments, investors, and so on. These data include, for example, population, employment or income per capita by geographic area. The outstanding characteristics of cross-sectional data are a great range of values and systematically varying error variances and coefficients of variation. Predictions encompass both forecasts and estimates, the only difference, generally irrelevant to this article, is the date being predicted relative to the date the predictions are made. In the context of predictions, accuracy is generally a subjective concept. [^1] Compare Lindley (1953, p. 46): “…I feel that loss is not easily separated in our minds from beliefs.” Later we develop loss functions to reflect beliefs. Even if objective measures of scale can be used, they may not necessarily coincide with the user’s concept of loss. We axiomatically develop loss functionns, which measure the “badness” of errors as a function of the sizes of the error and the actual values, or, equivalently, the predicted and actual values. This approach is in contrast to the common use of measures based on time series. [^2] Armstrong and Collopy (1992) and Fildes (1992), to give two examples, are quite clear about the time series basis of the error measures they use. These measures were developed to measure variations in the level of a single variable over time and are, thus, inapplicable to cross-sectional data. [^3] NRC (1980), is one of the few examples to use cross-sectional criteria: “‘To minimize errors in allocations’ is a vague statement which allows several interpretations (NRC, 1980, p.86).” The allocations are revenues directed to states and localities in proportion to their populations. I am indebted to Bashiruddin Ahmed for providing this example. Spencer (1980, 1985, 1986) also uses cross-sectional criteria. Initially, an impartial decision-maker is assumed to exist and asked to specify values of his loss function (disguised) for various combinations of errors and actual values. The decision-maker is assumed to be interested solely in the overall accuracy of the predictions and not to have an interest in the accuracy for a particular area or set of areas. The context of the predictions is also assumed fixed. [^4] I’d like to thank Stan Smith for pointing this out. Linear regression is then used to estimate the parameters of the loss function. [^5] The assumption of an impartial decision-maker and the statistical construction of the loss function provide an answer Spencer’s (1986) criticisms about whose preferences are to be incorporated into a von Neumann-Morgenstern loss function and how. These constructions are very much along the lines of Wald (1950) and Lindley (1953, 1985). An upshot of this process is that, in general, no single “ideal” measure of accuracy exists, as the evaluation of accuracy depends on the evaluator. [^6] Tukey (1986, p. 408) writes in the context of cross-sectional population estimates: “‘Optimum’ estimates, for example, are only optimum under narrow specifications that do not hold exactly in practice.” (Original underline) In the special case in which predictions represent resource allocations, the apportionment literature is used to establish the Webster-Saint Lagüe Rule as the appropriate loss function.
@@
Section 2 axiomatically constructs the simple loss function that we use. It further specifies assumptions and develops one of the simplest loss functions that satisfy those assumptions and has the property that the loss associated with a given absolute relative error [^7] Equivalently, one can also refer to the absolute percentage error (APE), but this will complicate the arguments made later on. APE will appear, however, in the example in Section 6. increases in the actual value. Section 3 discusses estimation of this loss function, along with specification tests. Different sets of predictions may have systematic, offsetting biases. The decision-maker may wish to form weighted averages of these predictions to produce a set of new predictions. Subection 3.4 provides a method to search for the optimal weightings which minimize the total loss function. Section 4 argues for the use of the Webster-Saint Lagüe Rule when predictions can be viewed as apportionments. Given a set of predicted and actual values, the total loss function is the sum of the loss functions for the individual observations. Section 5 provides an example of the use of loss functions, comparing the Webster-Saint Lagüe Rule to the commonly used mean absolute percentage error (MAPE).
@@ 2 The Loss Function
This section describes the assumptions used to generate the loss function, $L(P;A)$, where $P$ is the predicted value and $A$ is the actual value. [^8] Muhsam (1966) was apparently the first to use this form of loss in the context of predicting population. He was concerned with the loss associated with a point forecast of the population of a single geographic area. He assumed continuity of loss and a version of Assumption 2 below. I’d like to thank Jay Siegel for this point. After the assumptions are made, onr of the simplest forms of $L$ which satisfies them is specified. Restrictions on the values of the parameters of $L$ which make it increase in $A$ for a given relative error are then specified. The total loss $\mathcal{L}$ is the sum of the losses associated with each observation $i$: $\mathcal{L}=\sum_{i}^{n}L(P_{i},A_{i})$, where $n$ is the number of observations. [^9] In terms of Spencer (1986), $L$ is a component loss function and $\mathcal{L}$ is the loss function. The U.S. Census Bureau also uses this terminology for evaluating census adjustments. While other forms of $\mathcal{L}$ are possible, the summation corresponds to the concept of additively separable utility and obeys the von Neumann-Morgenstern expected utility axioms. [^10] That is, loss is the negative of utility. Additive losses are equivalent to additive utility. The von Neumann-Morgenstern expected utility axioms imply that the expected utility of a gamble is equal to the utilities of the outcomes multiplied by their probabilities. In terms of loss, the expected loss when losses of values $L_{1}$ and $L_{2}$ occur with probabilities $p_{1}$ and $p_{2}$, respectively, is $p_{1}L_{1}+p_{2}L_{2}$. This approach is similar to Lindley’s (1985, p. 59) commandment to “Go forth and maximise your expected utility.” (Original quotes) Lindley (1985) builds his book around the concept of making decisions that maximize a decision-maker’s expected utility. Since expectation requires specifying probability distributions, which we do not do, our commandment can be stated as “Go forth and maximize utility.”
@@
One of the simplest function swhich satisfies Assumptions 1–3 and admits Property 1 below is the Cobb-Douglas function
@@ 3.2 Determining the Loss Function at a Set of Points and Estimating it Globally
where $L_{j}>0,a$ is the constant of integration, an ignorable scaling factor, and $u_{j}$ is the regression residual. The values of $p$ and $q$ found from this regression become the parameters of the loss function. One caveat is in order: if the decision-maker reports a $U_{j}=100$, then $L_{j}=0$ and $\log L_{j}=-\infty$. The fix for this is to either delete this observation from the analysis or to choose a small $\delta>0$ to bound the $L_{j}$ from below.
@@ 3.2.1 Multiple Decision Makers
A final alternative is to adopt a loss function which works “well” in most cases as a convention in the manner of Keyfitz (1979). The Mean Absolute Percentage Error (MAPE,) discussed in Section 6, is frequently used in this sense. Another alternative is the Webster-Saint Lagüe Rule, discussed in Section 5. The Webster-Saint Lagüe Rule is a particular parametrization of our loss functions. That the chosen loss function need not work well in every case has been explicated by Lerner (1957, p. 441) in the context of legal rules:
@@ 3.3 Specification Tests
While equations (1a) and (1b) are the simplest form that thid loss function can take, they are hardly the only forms. In fact, an infinite number of forms of loss functions are possible. One quick specification test is to look at the sum $p+q$. If it is nonpositive, then it violates Property 1. Either the equation is misspecified, [^23] Remember that equations (1a) and (1b) represent only one of an infinite number of loss functions that satisfy Assumptions 1-3 and Property 1. or the decision-maker’s preferences violate Property 1.242424In some contexts, this may be reasonable. The form of the function may be misspecified, for which many tests are available. Heteroscedasticity may be present, which can be accounted for by weighted regression.
@@ 3.4 Optimal Averages of Predictions
Given two or more sets of predictions, it may be possible to improve on their accuracy by finding an optimal weighted average of these predictions. Suppose there are two sets of estimates $\mathbf{P}^{1}$ and $\mathbf{P}^{2}$, with associated total losses $\mathcal{L}^{1}$ and $\mathcal{L}^{2}$. Let $w_{1}$ and $w_{2}$ be weights such that $w_{1}+w_{2}=1$. By varying the values of $w_{1}$ and $w_{2}$, one creates new sets of predictions, one per set of weights. [^25] The new estimates may have to be constrained to sum up to a predetermined overall total or several predetermined subtotals. The total loss minimizing set of estimates and weights is then chosen. The optimal search method is unclear: this probably requires grid-searching. With $k>2$ estimate sets, this becomes a possibly expensive mesh search in $k-1$ dimensions.
@@ 4 Predictions as Apportionments: The Webster-Saint Lagüe Rule
Often predictions represent apportionments, that is, resource or legislative seat allocations. The U.S. decennial census is used to apportion the House of Representatives. Population estimates are used by some states to apportion tax revenues to localities. One would like to have the rule used for evaluating the predictions satisfy various fairness criteria. Balinski and Young (2001) show that the Webster-Saint Lagüe Rule satisfies a large number of fairness criteria. [^29] However, while Arrow’s(1950) Impossibility Theorem implies that no voting system can simultaneously satisfy all fairness criteria, the Webster-Saint Lagüe Rule is the only apportionment rule considered by Balinski and Young (1979, 2001) and Ernst (1992) that can be represented by minimizing the sum of loss functions satisifying Assumptions 1-3 and Property 1 (Spencer, 1985). I would like to thank Gregg Diffendal for informing me of Ernst (1992). Most of these criteria result from the Webster-Saint Lagüe Rule being a divisor method. The Webster-Saint Lagüe Rule is the only divisor method considered by Balikski and Young which has an admissible (in the present paper) loss function representation: $p=2$ and $q=-1$, (Balinski and Young, 2001) See Subsection 6.5 below for other apportionment-based loss functions. An additional normative criterion of particular interest is unbiasedness: for any two disjoint groups of areas, with all members in one group larger than those in the other, the probability that the method favors large areas equals the probability that it favors the small areas. (Balinksi and Young, 2001, Theorem 5.3)
@@ 4.1 An A Priori Probabilistic Interpretation of the Webster-Saint Lagüe Rule
$=\mathrm{Var}\left(\sum_{k=1}^{A_{i}}\eta_{i}^{j}\right)$ (6d)
$=A_{i}\mathrm{Var}(\eta_{i}^{j})$ (6e)
@@
where equation (6d) follows from asssumed unbiasedness, equation (6e) follows from the variance of the sum of independent random variables being equal to the sum of their variances (equal in this example) and equation (6f) follows from the definition of $\sigma^{2}_{j}$. Therefore,
@@
Thus, the Webster-Saint Lagüe rule produces estimates of the variances of each observation, $A_{i}\hat{\sigma}_{ij}^{2}$. Using its minimization to select predictions chooses the estimates set with the lowest $\hat{\sigma}_{j}^{2}$. Relaxing the the identical variance assumption so that $\sigma^{2}_{ij}$ is not constant causes this rule to estimate the average variance $\hat{\bar{\sigma}}^{2}_{j}=\frac{1}{n}\sum_{k=1}^{n}\hat{\sigma}_{kj}^{2}$. Now, the Webster-Saint Lagüe loss minimization selects the predictions set with the lowest average estimated variance. This is robust to small departures from the assumptions.
@@ 5 Example of Evaluating Predictions Using Loss Functions
Three Scenarios are used to compare the results of an evaluation using a loss function to those obtained by using MAPE. Scenario 1 is the baseline scenario with $\mathrm{APE}_{i})\equiv 2$ and $\mathcal{L}_{W}/n=11.08$. In Scenario 2, $\mathrm{APE}_{i}$ is reduced to 1 for $i\leq 5$, but $\mathrm{APE}_{6}$ increases to 10. That is, all but the smallest areas have their APEs halved, but the very smallest area’s APE increases by a factor of 5. MAPE increases to 2.5, but $\mathcal{L}_{W}/n$ falls to 2.9. Thus, MAPE ranks Scenario 2 as being less accurate than Scenario 1, even though the individual errors are smaller except for the very smallest area. On the other hand, using a loss function takes into account the size of the smallest area and discounts its accuracy loss and considers Scenario 2 to be more accurate. In Scenario 3, $\mathrm{APE}_{i}$ falls by 15% to 1.7 for $i=2\leq i\leq 5$, rises by 50% to 3 in the largest area, and is unchanged in the smallest area. MAPE falls slightly to 1.97, but $\mathcal{L}_{W}/n$ rises to 18.19. Thus, MAPE considers Scenario 3 to be superior to Scenario 1, as a result of the general reduction in the $\mathrm{APE}_{i}$, in spite of the major loss in accuracy in the largest area. The Webster-Saint Lagüe Rule loss function, on the other hand, puts a large weight on the accuracy loss in area 1 and increases its error measure relative to Scenario 1. Thus, this loss function puts increasing weight on an error as the size of the area increases. Putting all of these together, we find that MAPE and the Webster-Saint Lagüe loss function produce exactly opposite rankings of the Scenarios.